### v1

@@ Abstract
The quest for interpretable reinforcement learning is a grand challenge for the deployment of autonomous decision-making systems in safety-critical applications. Modern deep reinforcement learning approaches, while powerful, tend to produce opaque policies that compromise verification, reduce transparency, and impede human oversight. To address this, we introduce BASIL (Best-Action Symbolic Interpretable Learning), a systematic approach for generating symbolic, rule-based policies via online evolutionary search with quality-diversity (QD) optimization. BASIL represents policies as ordered lists of symbolic predicates over state variables, ensuring full interpretability and tractable policy complexity. By using a QD archive, the methodology in the proposed study encourages behavioral and structural diversity between top-performing solutions, while a complexity-aware fitness encourages the synthesis of compact representations. The evolutionary system supports the use of exact constraints for rule count and system adaptability for balancing transparency with expressiveness. Empirical comparisons with three benchmark tasks—CartPole-v1, MountainCar-v0, and Acrobot-v1—show that BASIL consistently synthesizes interpretable controllers with compact representations comparable to deep reinforcement learning baselines. Herein, this article introduces a new interpretable policy synthesis method that combines symbolic expressiveness, evolutionary diversity, and online learning through a unifying framework.
@@ 1 Introduction
Reinforcement Learning (RL) has proved a core methodology for solving complex decision-making and control problems in a range of domains, such as robotic control, autonomous vehicles, financial trading, and healthcare systems [1, 2]. Latest advancements in Deep Reinforcement Learning (DRL) techniques, based on strategies such as Deep Q-Networks (DQN)[3], Proximal Policy Optimization (PPO)[4], and Soft Actor-Critic (SAC)[5], have proved capable of achieving state-of-the-art performance for tasks with high-dimensional state and action spaces. Nonetheless, in spite of their empirical success, DRL algorithms tend to produce non-transparent, opaque policies, with important interpretability and reliability drawbacks in deployment domains wherein safety is a top concern. Transparency issues with such policies weaken the basis for building confidence and accountability, as well as the potential for human monitoring, in domains wherein transparent decision-making procedures are crucially important[6, 7].
@@
Additionally, the use of an advanced QD archive greatly enhances the evolutionary exploration process. By taking advantage of features like policy length and mean thresholds of predicates, our archive systematically guarantees complete coverage of behavior [15]. This new combination encourages the evolution of innovative, yet consistently high-performing strategies, thus enabling stakeholders to select from a range of policies crafted for a variety of operational environments or interpretability requirements. Therefore, our system not only successfully addresses the interpretability aspect but also greatly augments exploration in the solution space, generating a rich and diverse range of behaviors that are commonly ignored by traditional evolutionary approaches [13].
We demonstrate the effectiveness and interpretability of our approach through systematic experimental validation on standard benchmark reinforcement learning domains, such as CartPole, MountainCar, and Acrobot. Our experiments suggest our methodology generates a wide range of interpretable policies with state-of-the-art performance in comparison with top deep reinforcement learning techniques, while vastly outperforming traditional ERL-based approaches in terms of interpretability as well as policy diversity [16]. Our experimental methodology consists of exhaustive ablation evaluations, effectively demonstrating the effectiveness and useful contribution of every component in our proposed methodology. Further analysis discovers our quality diversity archive helps in the selection of a wide range of insightful policy variations, leading towards an advanced understanding of the underlying decision-making dynamics as well as emphasizing the real-world and theoretical advantages of our approach.
This paper introduces BASIL, a novel reinforcement learning framework for the construction of concise, symbolic policies through the evolution of genetic rules under the assistance of a Quality-Diversity archive. BASIL builds on the state-of-the-art in interpretable reinforcement learning by combining the online evolution of policies, a logical rule-based system, and diversity-focused optimization in a single framework. With experiments spanning a range of control benchmarks, we show that BASIL generates highly interpretable controllers with performance matching deep reinforcement learning and prior evolutionary state-of-the-arts in most cases. Our analysis of the diversity of policies in the Quality-Diversity archive highlights its essential role in driving the exploration of structurally diverse and semantically expressive policies and provides new insights into the interplay between exploration, expressiveness, and the interpretability of symbolic policy learning. As a result, BASIL constitutes a major innovation in the area of interpretable reinforcement learning, providing unparalleled transparency, flexibility, and range of solutions for complex decision-making tasks.
@@ 2 Related Work
The pursuit of interpretable reinforcement learning (RL) has garnered significant attention, leading to a myriad of approaches that attempt to balance performance and interpretability. Our approach stands out in that it evolves compact, symbolic rule-based policies through online environment interactions, with a particular focus on both interpretability and diversity. In what follows, we place our work in the context of existing research.
@@ 2.2 Post-hoc Interpretability and Program Synthesis
Programmatically Interpretable RL. Verma et al. proposed the idea of Programmatically Interpretable Reinforcement Learning (PIRL), which represents policies as programs in a high-level language. Within this approach, a neural policy acts as an oracle for generating interpretable programs with similar behavior as the neural policy [19]. While the policies learned by PIRL are interpretable and verifiable, the approach is post-hoc in nature and does not evolve policies directly through environment interactions.
@@ 2.4 Other Notable Approaches
Behavior Tree Extraction. Recent studies have explored deriving interpretable behavior trees from RL policies. These methods focus on translating learned policies into behavior trees, facilitating human understanding and verification. However, they often lack mechanisms for ensuring diversity or simplicity in the resulting structures.
Fuzzy Controllers and Neuro-Fuzzy Systems. Hein et al. also investigated the use of fuzzy controllers for interpretable RL, employing genetic programming and particle swarm optimization to evolve fuzzy rule sets [21]. While these approaches yield interpretable policies, they may require large rule sets, potentially compromising simplicity.
Symbolic Regression for Value Functions. Babuška discussed the application of genetic programming and symbolic regression to construct analytic value functions in RL [22]. This work focuses on value function approximation rather than direct policy derivation, offering insights into interpretable function approximation techniques.
Program Synthesis for Policy Learning. Liventsev et al. proposed BF++, a language designed for general-purpose program synthesis in RL contexts [23]. Their approach emphasizes the synthesis of interpretable programs representing policies, although it does not explicitly address policy diversity or simplicity.
Neuro-Fuzzy System Distillation. Gevaert et al. presented a method for distilling deep RL models into interpretable neuro-fuzzy systems [24]. This technique aims to combine the performance of deep RL with the interpretability of fuzzy systems, though it may not guarantee compact policy representations.
Symbolic Policy Learning. Landajuela et al. introduced Deep Symbolic Policy (DSP), a framework that searches for tractable mathematical expressions representing policies [25]. DSP emphasizes the discovery of symbolic policies with strong generalization capabilities, aligning with the goals of interpretability and simplicity.
Explainable RL via Symbolic Logic. Recent advancements have explored integrating symbolic logic into RL to enhance interpretability. These methods aim to produce explanations where key concepts and their relationships are described via intuitive symbols and rules, facilitating human understanding of agent behaviors.
@@ 3.1 Policy Representation
  • denotes a logical operator, specifically either less-than () or greater-than (),
@@
Furthermore, the action for every rule is chosen from a small, predefined set of possible actions defined in the structure of the environment. The final piece of the policy representation, the fallback rule, supports resilience by specifying a default action to be taken in the event that none of the conditions of the relevant rules are satisfied. Typically, the fallback action defaults to a conservative or safe alternative (typically specified as action 0), thus ensuring the policy remains unambiguously defined and executable in every environmental context.
One of the essential features of our approach to representing policies is the built-in ability to determine and limit the overall number of rules. By specifying the desired number of rules explicitly, users and domain experts are given precise control over the level of complexity in the policy. This feature enables fine-tuning based on practical factors such as computational resources, interpretability requirements, or specific demands of the application domain. Our methodology, therefore, not only supports the discovery of such policies but also ensures they meet the interpretability and operational requirements relevant to their deployment context.
Finally, the structured, symbolic rule-based policy representation provides policies with clear, understandable, and interpretable directives. The deliberate and systematic design of condition-action rules significantly enhances transparency and facilitates human validation and verification, making it especially suitable for use in safety-related and high-consequence decision-making applications.
@@ 3.2 Genetic Operators
The evolution cycle starts with an initialization step, wherein an initial population of policies is randomly generated. Each policy consists of a fixed or random number of rules, with every rule containing logical predicates over the state variables, paired with a corresponding discrete action. By this methodology, an initial diversity is ensured, thus creating a complete and representative starting point for further exploration.
@@ 5 Conclusion
We are not the first to argue that interpretable models matter. But with BASIL, we go further: we show that interpretability can be *evolved*, not engineered; *inherent*, not imposed; and *essential*, not optional. In doing so, we provide not just a tool for researchers, but a blueprint for a new generation of transparent learning systems—systems that are fast, capable, and always explainable.
Final Thoughts. The future of reinforcement learning does not lie in larger black boxes. It lies in clarity. It lies in methods like BASIL, where every decision has a reason, every action has a logic, and every policy is a story we can read. As we move toward human-aligned AI, symbolic policy learning may not just be an alternative—it may be the path we need.

### v3

@@ Abstract
The quest for interpretable reinforcement learning is a grand challenge for the deployment of autonomous decision-making systems in safety-critical applications. Modern deep reinforcement learning approaches, while powerful, tend to produce opaque policies that compromise verification, reduce transparency, and impede human oversight. To tackle this, we introduce BASIL (Best-Action Symbolic Interpretable Learning), a systematic approach for generating symbolic, rule-based policies via online evolutionary search with quality-diversity (QD) optimization. BASIL represents policies as ordered lists of symbolic predicates over state variables, ensuring full interpretability and tractable policy complexity. By using a QD archive, the methodology in the proposed study encourages behavioral and structural diversity between top-performing solutions, while a complexity-aware fitness encourages the synthesis of compact representations. The evolutionary system supports the use of exact constraints for rule count and system adaptability for balancing transparency with expressiveness. Empirical comparisons with three benchmark tasks—CartPole-v1, MountainCar-v0, and Acrobot-v1—show that BASIL consistently synthesizes interpretable controllers with compact representations comparable to deep reinforcement learning baselines. Herein, this article introduces a new interpretable policy synthesis method that combines symbolic expressiveness, evolutionary diversity, and online learning through a unifying framework.
@@ 1 Introduction
Reinforcement Learning has demonstrated itself as a core method for solving complex decision-making and control problems in a range of domains, such as robotic control, autonomous vehicles, financial trading, and healthcare systems [1, 2]. Latest advancements in Deep Reinforcement Learning (DRL) techniques, based on strategies such as Deep Q-Networks (DQN)[3], Proximal Policy Optimization (PPO)[4], and Soft Actor-Critic (SAC)[5], have proved capable of achieving state-of-the-art performance for tasks with high-dimensional state and action spaces. Nonetheless, in spite of their empirical success, DRL algorithms tend to produce non-transparent, opaque policies, with important interpretability and reliability drawbacks in deployment domains wherein safety is a top concern. Transparency issues with such policies weaken the basis for building confidence and accountability, as well as the potential for human monitoring, in domains wherein transparent decision-making procedures are crucially important[6, 7].
@@
Additionally, the use of an advanced QD archive greatly enhances the evolutionary exploration process. By taking advantage of features like policy length and mean thresholds of predicates, our archive systematically guarantees complete coverage of behavior [15]. This new combination encourages the evolution of innovative, yet consistently high-performing strategies, thus enabling stakeholders to select from a range of policies crafted for a variety of operational environments or interpretability requirements. Therefore, BASIL not only successfully addresses the interpretability aspect but also greatly augments exploration in the solution space, generating a rich and diverse range of behaviors which are commonly ignored by traditional evolutionary approaches [13].
We show the effectiveness and interpretability of BASIL through systematic experimental validation on standard benchmark reinforcement learning domains, namely CartPole, MountainCar, and Acrobot. Our experiments suggest our methodology generates a wide range of interpretable policies with state-of-the-art performance in comparison with top deep reinforcement learning techniques, while vastly outperforming traditional ERL-based approaches in terms of interpretability as well as policy diversity [16]. Our experimental methodology consists of exhaustive ablation evaluations, effectively demonstrating the effectiveness and useful contribution of every component in our proposed methodology. Further analysis discovers our quality diversity archive helps in the selection of a wide range of insightful policy variations, leading towards an advanced understanding of the underlying decision-making dynamics as well as emphasizing the real-world and theoretical advantages of our approach.
This paper introduces BASIL, a novel reinforcement learning framework for the construction of concise, symbolic policies through the evolution of genetic rules under the assistance of a Quality-Diversity archive. BASIL builds on the state-of-the-art in interpretable reinforcement learning by combining the online evolution of policies, a logical rule-based system, and diversity-focused optimization in a single framework. With experiments spanning a range of control benchmarks, we show that BASIL generates highly interpretable controllers with performance matching deep reinforcement learning and prior evolutionary state-of-the-arts in most cases. Our analysis of the diversity of policies in the Quality-Diversity archive highlights its essential role in driving the exploration of structurally diverse and semantically expressive policies and provides new insights into the interplay between exploration, expressiveness, and the interpretability of symbolic policy learning. Accordingly, BASIL constitutes a major innovation in the area of interpretable RL, providing unparalleled transparency, flexibility, and range of solutions for complex decision-making tasks.
@@ 2 Related Work
The pursuit of interpretable reinforcement learning has caught significant attention, leading to a great variety of approaches that attempt to balance performance and interpretability. Our approach stands out in that it evolves compact, symbolic rule-based policies through online environment interactions, with a particular focus on both interpretability and diversity. In what follows, we place our work in the context of existing research.
@@ 2.2 Post-hoc Interpretability and Program Synthesis
Programmatically Interpretable RL. Verma et al. proposed the idea of Programmatically Interpretable Reinforcement Learning (PIRL), which represents policies as programs in a high-level language. Within this approach, a neural policy acts as an oracle for generating interpretable programs with similar behavior as the neural policy [19]. Although the policies learned by PIRL are interpretable and verifiable, the method is retrospective and does not learn policies actively through environment interactions.
@@ 2.4 Other Notable Approaches
Behavior Tree Extraction. Recent studies have explored deriving interpretable behavior trees from RL policies. These methods focus on translating learned policies into behavior trees, facilitating human understanding and verification. Nevertheless, they often lack mechanisms for ensuring diversity or simplicity in the resulting structures.
Fuzzy Controllers and Neuro-Fuzzy Systems. Hein et al. studied the use of fuzzy controllers for interpretable reinforcement learning, leveraging genetic programming and particle swarm optimization to evolve fuzzy rule sets [21]. While these methods result in interpretable policies, they may need large rule sets, potentially compromising simplicity.
Symbolic Regression for Value Functions. Babuška investigated the potential of genetic programming and symbolic regression to obtain symbolic expressions of value functions in RL [22]. This work concentrates on value function approximation rather than direct policy derivation, yielding insights into interpretable function approximation techniques.
Program Synthesis for Policy Learning. Liventsev et al. proposed BF++, a language designed for general-purpose program synthesis in RL contexts [23]. Their approach emphasizes the design of interpretable programs encoding policies, though it does not clearly account for policy diversity or simplicity.
Neuro-Fuzzy System Distillation. Gevaert et al. presented a method for distilling deep RL models into interpretable neuro-fuzzy systems [24]. This framework attempts to integrate the performance of deep reinforcement learning into the interpretability of fuzzy systems, though it may not guarantee compact policy representations.
Symbolic Policy Learning. Landajuela et al. introduced Deep Symbolic Policy (DSP), a method that investigates for tractable mathematical expressions representing policies [25]. DSP emphasizes the discovery of symbolic policies with strong generalization capabilities, aligning with the goals of interpretability and simplicity.
Explainable RL via Symbolic Logic. Recent studies have explored integrating symbolic logic into RL to enhance interpretability. These approaches aim to produce explanations where key concepts and their relationships are described via intuitive symbols and rules, enabling human understanding of agent behaviors.
@@ 3.1 Policy Representation
  • denotes a logical operator, specifically either less-than (¡) or greater-than (¿),
@@
Hence, the action for every rule is chosen from a small, predefined set of possible actions defined in the structure of the environment. The last piece of the policy representation, the fallback rule, supports resilience by specifying a default action to be taken in the event that none of the conditions of the relevant rules are satisfied. Typically, the fallback action defaults to a conservative or safe alternative (typically specified as action 0), thus ensuring the policy remains unambiguously defined and executable in every environmental context.
One of the main features of our approach to representing policies is the built-in ability to determine and limit the overall number of rules. By specifying the desired number of rules explicitly, users and domain experts are given precise control over the level of complexity in the policy. This feature enables fine-tuning based on practical factors for example computational resources, interpretability requirements, or specific demands of the application domain. Our methodology, therefore, not only supports the discovery of such policies but also ensures they meet the interpretability and operational requirements relevant to their deployment context.
Finally, the structured, symbolic rule-based policy representation provides policies with understandable and interpretable directives. The deliberate and systematic design of condition-action rules significantly enhances transparency and facilitates human validation and verification, making it especially suitable for use in safety-related and high-consequence decision-making applications.
@@ 3.2 Genetic Operators
The evolution cycle starts with an initialization step, wherein an initial population of policies is randomly generated. Each policy consists of a fixed or random number of rules, with every rule containing logical predicates over the state variables, paired with a corresponding discrete action. By this methodology, an initial diversity is ensured, hence creating a complete and representative starting point for further exploration.
@@ 5 Conclusion
We are not the first to argue that interpretable models matter. But with BASIL, we go further: we show that interpretability can be evolved, not engineered; inherent, not imposed; and essential, not optional. In doing so, we provide not just a tool for researchers, but a blueprint for a new generation of transparent learning systems—systems that are fast, capable, and always explainable.
The future of reinforcement learning does not lie in larger black boxes. It lies in clarity. It lies in methods like BASIL, where every decision has a reason, every action has a logic, and every policy is a story we can read. As we move toward human-aligned AI, symbolic policy learning may not just be an alternative—it may be the path we need.